OpenStudy (anonymous):

The implied domain of y = arcsin[cos(-2x)] is: A - [-pi/2 , 0] B - [0 , pi/2] C - [/pi , pi] D - [-pi/4 , pi/4] E [-2pi , 2pi]

OpenStudy (anonymous):

\[ \sin(y)=\cos(-2x) \]

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

\[ \cos(\pi/2-y) = \cos(-2x)\implies \pi/2-y = -2x \]

OpenStudy (anonymous):

I'm just messing around.

OpenStudy (anonymous):

Okay because I am really confused haha

OpenStudy (anonymous):

Okay so basically we want to know the minimum angles to get \[ \cos(-2x)=1,\;\cos(-2x)=-1 \]

OpenStudy (anonymous):

0 is one of them.

OpenStudy (anonymous):

Honestly, this question is really weird, isn't it?

OpenStudy (anonymous):

Yeah haha

OpenStudy (anonymous):

I don't see any real restriction on the domain.

OpenStudy (anonymous):

Oh well, don't worry, thanks anyway :)